find an equation of the line having the given slope and containing the given point.
m=6/7, (8,-1)

Answer by ewatrrr(24785)

 
Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
m=6/7, (8,-1)
y = (6/7)x + b
-1 = (6/7)·8 + b
-55/7 = b
y = (6/7)x -55/7
standard form of a line is in the form Ax + By = C: 6x - 7y = 55